# GPTKB: Comprehensively Materializing Factual LLM Knowledge

### v3

@@ Abstract
LLMs have majorly advanced NLP and AI, and next to their ability to perform a wide range of procedural tasks, a major success factor is their internalized factual knowledge. Since Petroni et al. (2019), analyzing this knowledge has gained attention. However, most approaches investigate one question at a time via modest-sized pre-defined samples, introducing an “availability bias” Tversky and Kahneman (1973) that prevents the discovery of knowledge (or beliefs) of LLMs beyond the experimenter’s predisposition.
To address this challenge, we propose a novel methodology to comprehensively materialize an LLM’s factual knowledge through recursive querying and result consolidation.
As a prototype, we employ GPT-4o-mini to construct GPTKB, a large-scale knowledge base (KB) comprising 101 million triples for over 2.9 million entities - achieved at 1% of the cost of previous KB projects. This work marks a milestone in two areas: For LLM research, for the first time, it provides constructive insights into the scope and structure of LLMs’ knowledge (or beliefs). For KB construction, it pioneers new pathways for the long-standing challenge of general-domain KB construction. GPTKB is accessible at https://gptkb.org.
GPTKB: Comprehensively Materializing Factual LLM Knowledge
Yujia Hu1   Tuan-Phong Nguyen2   Shrestha Ghosh3   Simon Razniewski1 1ScaDS.AI & TU Dresden, Germany 2Max Planck Institute for Informatics, Saarbrücken, Germany 3 University of Tübingen, Germany yujia.hu@mailbox.tu-dresden.de  {tuanphong,ghoshs}@mpi-inf.mpg.de  simon.razniewski@tu-dresden.de
@@ 1 Introduction
p0.44|p0.48 Existing LLM analyses Our proposal
@@
To enable comprehensive insights into LLM knowledge, we propose to extensively materialize the knowledge of LLMs into a KB (see Table 1). General-world KBs like Wikidata Vrandecic and Krötzsch (2014), Yago Suchanek et al. (2007) and DBpedia Auer et al. (2007) are important and long-standing backbones for AI applications, yet have seen little innovation in the last years, and have only been constructed manually, or by scraping semistructured web content Weikum et al. (2021). In particular, we propose to use recursive graph exploration to obtain comprehensive (named-entity-centric) LLM knowledge, and to consolidate it via LLM-based entity disambiguation, class and relation canonicalization, and taxonomy induction (see Fig. 1). This proposal faces several challenges:
@@
    Runtime and cost: State-of-the-art KBs contain millions of entries, and LLM inference is relatively slow. It is therefore unclear how to perform comprehensive knowledge elicitation under practical monetary and time constraints.
@@
    Global inconsistency: Distinctive prompts risks surfacing outputs that are not globally consistent, e.g., introducing duplicate relations, entities, or disconnected class hierarchies.
@@
    Using GPT-4o-mini, we construct GPTKB, the first large-scale KB entirely built from an LLM’s parametric knowledge, containing 101M assertions for over 2.9M entities.
@@
    We use GPTKB to exemplarily analyze GPT-4o-mini’s factual knowledge in terms of scale, accuracy, bias, and cutoff, at the same time.
Our work is a significant advancement for two areas: For LLM research, we introduce a proof-of-concept methodology that, for the first time, enables constructive insights into the knowledge (or beliefs) of LLMs. For the Semantic Web, we provide fresh momentum for the long stale problem of open-domain KB construction. We provide code and a concrete resource, GPTKB, both as a 3.8 GB download, and via an online browsing interface and SPARQL query interface at https://gptkb.org.
@@ 2 Related work
Since the emergence of LLMs, the question of how much these models know has frequently been raised Petroni et al. (2019); Roberts et al. (2020); Jiang et al. (2020); Veseli et al. (2023); Sun et al. (2024); Wu et al. (2024). So far, the widely adopted approach is to sample a domain of interest, e.g., question-answer pairs from a common benchmark, or triples from Wikidata, transiently probe the LLM, and compute the fraction that the LLM can correctly answer/complete. This is prone to “availability bias” Tversky and Kahneman (1973) since we do not get a comprehensive view of the LLM knowledge outside the focus of existing benchmarks. For instance, we found that GPT-4o-mini holds substantial knowledge about art periods or hobbies, which are not covered in existing KBs. Kassner et al. (2021) propose a small-scale persistent memory component for ensuring LLM answers remain consistent over multiple prompts. Nguyen et al. (2024) iteratively prompt GPT-3.5 for obtaining 167K sentences containing cultural commonsense for 11K subjects. Cohen et al. (2023) propose to iteratively prompt GPT-3 for relations and relational assertions for triples. In difference to Nguyen et al. (2024), we are after structured content (triples), at a much larger scale. In difference to Cohen et al. (2023), we unveil the practical challenges that such a graph exploration faces, tackle them, and perform LLM-based KB construction at scale.
@@
Iterative information extraction is a long-standing idea, already prototyped by Google cofounder Sergey Brin via the DIPRE system Brin (1998). The Snowball system by Agichtein and Gravano (2000) presented a large-scale implementation of this idea, harvesting tuples from over 300,000 newspaper documents. The KnowItAll system Etzioni et al. (2004) followed a similar approach, and led to the ReVerb KB Fader et al. (2011). Despite a proposal by Cohen et al. (2023), to date, we are not aware of any attempts to extract knowledge from LLMs at scale.
@@
Table 2: Size comparison of major KBs. See Appendix F for sources.
Large-scale KB construction
3 Methodology
@@ 3.1 Knowledge elicitation
The process starts from one or a set of seed subjects, e.g., Vannevar Bush, the visionary behind the concept of hyperlink-based knowledge organization Bush (1945). From the objects in the triples obtained for him, we can then identify further entities (e.g., MIT (affiliation) or Everett, MA (birth place)), for which we can then elicit further knowledge, and so on Brin (1998). Figure 2 illustrates how in 3 hops, we arrive at entities of diverse types, such as historical event (Boston Tea Party, Manhattan Project), newspaper (The Times), and magazine (The New Republic).
@@
Knowledge elicitation
A major challenge is to elicit as much knowledge as possible, but at the same time, not encourage hallucinations. We found that without guidelines on the expected number of triples, LLMs showed too little variance in the number of triples per subject, while we would expect them to return many more triples for Einstein than for other entities. We solve this via indications that are defined in relation to the entity’s popularity. In difference to the proposal Cohen et al. (2023), we also drop the separate relation elicitation, and relation-specific knowledge extraction, as these do not scale. To structure the knowledge, we also request that at least one instanceOf triple should be returned. Output parsing is eased by using the structured outputs feature of OpenAI’s API. In turn, this also reduces hallucinations, e.g., additional qualifiers, textual descriptions, or similar. The full prompt is in Figure 8 in the Appendix.
@@ 3.2 Knowledge consolidation
8:     $\textit{threshold}\leftarrow\alpha\times\frac{\log(\text{frequency}(r))}{\log(\text{frequency}(\text{first}(\mathcal{R})))}$
@@
Similarly to relations, there are 103K distinct class names (i.e., objects for instanceOf relations) generated by the knowledge elicitation phase, with both obvious duplicates, and many overly specific cases. Algorithm 1 is also applied to clean this set.
In our experiment, both relation and class clustering uses SentenceTransformers Reimers and Gurevych (2019) for embedding cosine similarity computation, with $\alpha=1.2$ chosen via manual inspection of small held-out sets.
@@
The classes in the KB so far do not form a coherent, or even connected taxonomy, as the knowledge elicitation process only expands named entities. We propose Algorithm 2, which is based on the LLM itself, to build a complete taxonomy from these individual classes.
@@
The prompts used for seed taxonomy generation, generality scoring, superclass identification, and taxonomy update are provided in Figure 10 in Appendix C. Further LLM-based refinement in the style of Peng et al. (2024) could be considered.
@@ 4 Implementation
Seed entity, result size, runtime and cost
@@
Our KB contains a total of 101M triples for 2.9M entities, organized into 1,804 relations and 473 classes. This gives an average of 36 triples per subject, with two distinct clusters, in particular, 651K entities with 10 triples and 86K with 50 triples, with most others near these values. Of all triples, more than 37M have an entity as object, and more than 67M have a literal as an object. The average length of entity labels is 24.5 characters. Example output for Vannevar Bush is shown in Figure 2. Statistics are shown in Table 3.
@@
Relations 1,804 (788K before canonicalization)
Classes 473 (103K before canonicalization)
@@
Avg. triples/entity 36
Avg. label length 24.5 characters
@@
We provide our KB both as a download (3.8 GB in TTL format), via a web browsing interface, and via a SPARQL query endpoint at https://gptkb.org. We provide our KB under the permissive CC BY-NC 4.0 license, though we caution against unreflected downstream usage (see Section 8). This license is compatible with the permissive terms of use of OpenAI OpenAI (2022).
@@ 5.1 Precision
For verifying entities, we retrieved 5 search results from a web search API for the entity label, which are then used to decide whether the entity’s existence could be verified, appears plausible, or could not be verified. Deciding on these labels based on textual context represents a task of textual entailment, a.k.a. natural language inference (NLI), a task generally considered solved for LLMs. [^2] On LLM performance on textual entailment, the popular SNLI Bowman et al. (2015) is comparable, which was considered solved in 2020 by BERT architectures Zhang et al. (2020) at 92% accuracy, which is the ceiling, given dataset noise. This is also why the literature mentions no numbers for newer models like Llama on this task anymore. We therefore use another LLM, Llama-3.1-70B-Instruct Dubey et al. (2024), for this task. The LLM hereby only judges textual entailment, it does not independently judge truth from its parameters. Based on 1,000 samples, average entity verifiability is 74%, and as shown in Figure 3, we observe a continuous drop of verifiability over the layers, stabilizing at layer 10 at about 70%. The plot also includes Wikidata-existence, which paints a similar trend at lower levels. We generally observe a wide variance across classes, finding, for instance, that persons consistently have a higher verifiability than fictional character, where many codes were made up (e.g., Officer K.I.T.T. XV and Officer K.I.T.T. XVI) or varying incorrect type numbers are added to real names continuously by the LLM (e.g., Agent 71 and Agent 72).
@@ 5.2 GPTKB content and comparison
The most frequent classes generated for entities are shown in Figure 4, where we find that Person dominates the dataset. Several more classes are actually merged into Person, e.g., human, hominid. Including these, the total number of Person is more than 287K. Further details by layer are in Appendix B.
@@
Figure 4: The 10 most frequent classes in GPTKB, which make up 25% of all entities.
Properties
The most frequent properties in the whole KB are instanceOf (3.1M) and features (2M) with similarly generic properties following at the entity level. At the class level, we find more specific properties, for instance, the most frequent properties in the class Person are instanceOf (309K), occupation (198K), knownFor (179K), exhibition (178K).
@@
We also analyze the geographic bias of our KB, using the listed nationalities as proxy. We observe a clear bias towards English-language nationalities: the top-2 are American (119K) and British (35K), while the next 3 are French (18K), German (14K), and Japanese (11K). This bias in GPTKB is stronger than, e.g., in Wikidata Shaik et al. (2021), and a curious lack of others (e.g., Chinese only 3K, compared to 8K for Russian, or 7K for Indian/Spanish), likely reflecting the (undisclosed) English-centric training corpus of GPT-4o-mini.
An interesting point is also gender bias: The gender property of GPTKB contains 15K female versus 8K male values. For first names, male ones are still more than female (47% vs. 37% based on the gender-guesser 0.4.0 Python library), but this is still a much lower bias than in other KBs, likely reflecting the effects of gender debiasing.
@@
We compare with Wikidata on several aspects: First, we compute the fraction of subjects that exist in Wikidata. From a random sample of 2,000 subjects from GPTKB, 24% have an entity with exactly matching label in Wikidata. A further 6.5% have a non-empty search result, i.e., an entity of paraphrased or similar label. The remaining 69.5% appear novel. Second, we exemplarily look at the 41 triples for Vannevar Bush, of which we find that more than 10 are not contained in Wikidata, e.g., his affiliation with the US government, his children count (incorrect by 1), or him inventing the concept of hypertext. [^3] An exact comparison is not straightforward, because relation names do not perfectly align, and objects in GPTKB are substantially more wordy. Third, we identify several properties not modelled at all in Wikidata, for instance, historicalSignificance (396K triples), artStyle (174K triples), or hobbies (44K triples). This indicates that GPTKB potentially contains a significant amount of novel knowledge. A more comprehensive KB comparison in the style of Färber et al. (2018) is planned.

### v4

# Enabling LLM Knowledge Analysis via Extensive Materialization

@@ Abstract
Large language models (LLMs) have majorly advanced NLP and AI, and next to their ability to perform a wide range of procedural tasks, a major success factor is their internalized factual knowledge. Since Petroni et al. (2019), analyzing this knowledge has gained attention. However, most approaches investigate one question at a time via modest-sized pre-defined samples, introducing an “availability bias” Tversky and Kahneman (1973) that prevents the analysis of knowledge (or beliefs) of LLMs beyond the experimenter’s predisposition.
To address this challenge, we propose a novel methodology to comprehensively materialize an LLM’s factual knowledge through recursive querying and result consolidation. Our approach is a milestone for LLM research, for the first time providing constructive insights into the scope and structure of LLM knowledge (or beliefs).
As a prototype, we build GPTKB, a knowledge base (KB) comprising 101 million relational triples for over 2.9 million entities from GPT-4o-mini. We use GPTKB to exemplarily analyze GPT-4o-mini’s factual knowledge in terms of scale, accuracy, bias, cutoff and consistency, at the same time. GPTKB is accessible at
https://gptkb.org.
Enabling LLM Knowledge Analysis via Extensive Materialization
Yujia Hu1   Tuan-Phong Nguyen2   Shrestha Ghosh3   Simon Razniewski1 1ScaDS.AI & TU Dresden, Germany 2Max Planck Institute for Informatics, Saarbrücken, Germany 3 University of Tübingen, Germany {yujia.hu,simon.razniewski}@tu-dresden.de  tuanphong@mpi-inf.mpg.de  shrestha.ghosh@uni-tuebingen.de
@@ 1 Introduction
@p0.44|p0.48@ Existing LLM analyses Our proposal
@@
To enable comprehensive insights into LLM knowledge, we propose to extensively materialize the knowledge of LLMs into a KB (see Table 1). General-world KBs like Wikidata Vrandecic and Krötzsch (2014), Yago Suchanek et al. (2007) and DBpedia Auer et al. (2007) are important and long-standing backbones for AI applications, yet have seen little innovation in the last years, and have only been constructed manually, or by scraping semistructured web content Weikum et al. (2021). In particular, we propose to use recursive graph exploration to obtain comprehensive (named-entity-centric) LLM knowledge, and to consolidate it via LLM-based entity disambiguation, class and relation canonicalization, and taxonomy induction (see Figure 1). This proposal faces several challenges:
@@
    Runtime and cost: State-of-the-art LLMs potentially store millions of facts, and LLM inference is relatively slow. It is therefore unclear how to perform comprehensive knowledge elicitation under practical monetary and time constraints.
@@
    Global inconsistency: Consecutive prompts risks surfacing outputs that are not globally consistent, e.g., introducing duplicate relations, entities, or disconnected class hierarchies.
@@
    Using GPT-4o-mini, we construct GPTKB, the first large-scale materialization of LLM knowledge, comprising 101M assertions for over 2.9M entities.
@@
    We use GPTKB to exemplarily analyze GPT-4o-mini’s factual knowledge in terms of scale, accuracy, bias, cutoff, and consistency, at the same time.
We provide code and a concrete resource, GPTKB, both as a 3.8 GB download, and via an online browsing interface and SPARQL query interface at https://gptkb.org.
@@ 2 Related work
Since the emergence of LLMs, the question of how much these models know has frequently been raised Petroni et al. (2019); Roberts et al. (2020); Jiang et al. (2020); Singhania et al. (2022); Veseli et al. (2023); Sun et al. (2024); Wu et al. (2025). So far, the widely adopted approach is to sample a domain of interest, e.g., question-answer pairs from a common benchmark, or triples from Wikidata, transiently probe the LLM, and compute the fraction that the LLM can correctly answer/complete. This is prone to “availability bias” Tversky and Kahneman (1973) since we do not get a comprehensive view of the LLM knowledge outside the focus of existing benchmarks. For instance, we found that GPT-4o-mini holds substantial knowledge about art periods or hobbies, which are not covered in existing KBs. Kassner et al. (2021) propose a small-scale persistent memory component for ensuring LLM answers remain consistent over multiple prompts. Nguyen et al. (2024) iteratively prompt GPT-3.5 for obtaining 167K sentences containing cultural commonsense for 11K subjects. Cohen et al. (2023) propose to iteratively prompt GPT-3 for relations and relational assertions for triples. In difference to Nguyen et al. (2024), we are after structured content (triples), at a much larger scale. In difference to Cohen et al. (2023), we unveil the practical challenges that such a graph exploration faces, tackle them, and perform LLM-based KB construction at scale.
@@
Iterative information extraction is a long-standing idea, already prototyped by Google cofounder Sergey Brin via the DIPRE system Brin (1998). The Snowball system by Agichtein and Gravano (2000) presented a large-scale implementation of this idea, harvesting tuples from over 300,000 newspaper documents. The KnowItAll system Etzioni et al. (2004) followed a similar approach, and led to the ReVerb KB Fader et al. (2011). Several recent works attempt IE with LLMs from text Hao et al. (2022); Carta et al. (2023); Zhang et al. (2024), but, except of the LM-KBC challenge series Singhania et al. (2022) and a proposal by Cohen et al. (2023), to date, we are not aware of any attempts to extract knowledge from LLMs.
Large-scale KB construction
@@
Table 2: Size comparison of major KBs. Sources in Appendix D.
3 LLM knowledge materialization
@@ 3.1 Knowledge elicitation
The process starts from one or a set of seed subjects, e.g., Vannevar Bush, the visionary behind the concept of hyperlink-based knowledge organization Bush (1945). We prompt the LLM to return knowledge about this subject in the form of triples. From the objects in the triples obtained for him, we then identify further entities using NER (e.g., MIT (affiliation) or Everett, MA (birth place)). Thus, we iteratively expand the entity graph and elicit further knowledge on the newly encountered entities Brin (1998). We elaborate on the knowledge prompting and NER in the following paragraphs. Figure 2 illustrates how in 3 hops, we arrive at entities of diverse types, such as historical event (Boston Tea Party, Manhattan Project), newspaper (The Times), and magazine (The New Republic).
@@
Knowledge prompting
A major challenge is to elicit as much knowledge as possible, but at the same time, not encourage hallucinations. We found that without guidelines on the expected number of triples, LLMs showed too little variance in the number of triples per subject, while we would expect them to return many more triples for Einstein than for other entities. We solve this via indications that are defined in relation to the entity’s popularity. In difference to the proposal Cohen et al. (2023), we also drop the separate relation elicitation, and relation-specific knowledge extraction, as these do not scale. To structure the knowledge, we also request that at least one instance_of triple should be returned. Output parsing is eased by using the structured outputs feature of OpenAI’s API. In turn, this also reduces hallucinations, e.g., additional qualifiers, textual descriptions, or similar. The full prompt is in Figure 9 in the Appendix.
@@ 3.2 Knowledge consolidation
8:     $\textit{threshold}\leftarrow\text{clip}\{\alpha\times\frac{\log(\text{frequency}(r))}{\log(\text{frequency}(\text{first}(\mathcal{R})))},\mathcal{H},\mathcal{L}\}$
@@
Similarly to relations, there are 103K distinct class names (i.e., objects for instance_of relations) generated by the knowledge elicitation phase, with both obvious duplicates, and many overly specific cases. Algorithm 1 is also applied to clean this set.
In our experiment, both relation and class clustering uses SentenceTransformers Reimers and Gurevych (2019) for embedding cosine similarity computation, with $\alpha=1.4$ and global highest threshold $\mathcal{H}=0.95$ chosen via manual inspection of small held-out sets. For tail predicates, to ensure clustering precision, we set lowest threshold $\mathcal{L}=0.75$.
@@
The classes in the KB so far do not form a coherent, or even connected taxonomy, as the knowledge elicitation process only expands named entities. But we also cannot simply prompt for a complete taxonomy to put on top (e.g., following the approach by Funk et al. (2023)), because we need to include the existing classes. We propose Algorithm 2, which is based on the LLM itself, to build a complete taxonomy that incorporates given classes.
@@
The prompts used for seed taxonomy generation, generality scoring, superclass identification, and taxonomy update are provided in Figure 11 in Appendix B. Further LLM-based refinement in the style of Peng et al. (2024) could be considered.
@@ 4 Implementation
Seed entity, runtime and cost
@@
Our KB contains a total of 101M triples for 2.9M entities, organized into 567K relations and 4,715 classes. This gives an average of 35 triples per subject, with two distinct clusters, in particular, 860K entities with 10 triples and 66K with 50 triples, with most others near these values. Of all triples, more than 37M have an entity as object, and more than 64M have a literal as an object. The average length of entity labels is 21.8 characters. Example output for Vannevar Bush is shown in Figure 2. Statistics are shown in Table 3.
@@
Relations 567K (788K before canonicalization)
Classes 4,715 (103K before canonicalization)
@@
Avg. triples/entity 35
Avg. label length 21.8 characters
@@
We provide our KB as a download (3.8 GB in TTL format), via a web browsing interface, and via a SPARQL query endpoint at https://gptkb.org under the permissive CC BY-NC 4.0 license, though we caution against unreflected downstream usage (see Section 8). This license is compatible with the permissive terms of use of OpenAI OpenAI (2022).
@@ 5.1 Precision
For verifying entities, we retrieved 5 search results from a web search API for the entity label, which are then used to decide whether the entity’s existence could be verified, appears plausible, or could not be verified. Deciding on these labels based on textual context represents a task of textual entailment, a.k.a. natural language inference (NLI), a task generally considered solved for LLMs. [^2] On LLM performance on textual entailment, the popular SNLI Bowman et al. (2015) is comparable, which was considered solved in 2020 by BERT architectures Zhang et al. (2020) at 92% accuracy, which is the ceiling, given dataset noise. This is also why the literature mentions no numbers for newer models like Llama on this task anymore. We therefore use another LLM, Llama-3.1-70B-Instruct Dubey et al. (2024), for this task. The LLM hereby only judges textual entailment, it does not independently judge truth from its parameters. Based on 1,000 samples, average entity verifiability is 74%, and as shown in Figure 3, we observe a continuous drop of verifiability over the layers, stabilizing at layer 10 at about 70%. The plot also includes Wikidata-existence, which paints a similar trend at lower levels. We generally observe a wide variance across classes, finding, for instance, that persons consistently have a higher verifiability than fictional characters, where many codes were made up (e.g., Officer K.I.T.T. XV and Officer K.I.T.T. XVI) or varying incorrect type numbers are added to real names continuously by the LLM (e.g., Agent 71 and Agent 72).
@@
[…]
Figure 4: The 10 most frequent classes in GPTKB, which make up 21.5% of all entities.
@@ 5.2 GPTKB content and comparison
The most frequent classes generated for entities are shown in Figure 4, where we find that the class Person dominates the dataset with 281K instances. This is followed by the classes film (66K) and Company (56K), which are are an order of magnitude smaller. Classes denoting creative works, such as, video game (46K), album (43K) and song (30K), also make it to the top ten frequent classes. Figure 5 shows the most frequent classes by BFS-level. As one can see, diversity increases towards lower BFS-levels, and some classes, like awards, only emerge later.
Properties
The most frequent properties in the whole KB are instance_of (3.1M) and features (2M) with similarly generic properties following at the entity level (see Figure 6(a)). At the class level, we find more specific properties, for instance, the most frequent properties in the class Person are instance_of (299K), occupation (191K), known_for (176K), nationality (166K), and in the class film, the properties character (399K), starring (134K) and genre (99K) are the most frequent, shown in Figures 6(b) and 6(c), respectively.
@@
(a) GPTKB.
[…]
(b) Class Person.
[…]
(c) Class film.
Figure 6: Distribution of 10 most frequent properties in GPTKB, and within classes Person and film.
@@
We also analyze the geographic bias of our KB, using the listed nationalities as proxy. We observe a clear bias towards English-language nationalities: the top-2 are American (119K) and British (35K), while the next 3 are French (18K), German (14K), and Japanese (11K), and Chinese and Indian only at 3K and 7K. This bias in GPTKB is stronger than, e.g., in Wikidata Shaik et al. (2021), likely reflecting the (undisclosed) English-centric training corpus of GPT-4o-mini.
An interesting point is also gender bias: The gender property of GPTKB contains 15K female versus 8K male values. For first names, male ones are still more than female (47% vs. 37% based on the gender-guesser 0.4.0 Python library), but this is still a much lower bias than in other KBs, likely reflecting efforts of gender debiasing the LLM.
@@
We compare with Wikidata Vrandecic and Krötzsch (2014) on several aspects: First, we compute the fraction of subjects that exist in Wikidata. From a random sample of 2,000 subjects from GPTKB, 24% have an entity with exactly matching label in Wikidata. A further 6.5% have a non-empty search result, i.e., an entity of paraphrased or similar label. The remaining 69.5% appear novel. Second, we exemplarily look at the 41 triples for Vannevar Bush, of which we find that more than 10 are not contained in Wikidata, e.g., his affiliation with the US government, his children count (incorrect by 1), or him inventing the concept of hypertext. [^3] An exact comparison is not straightforward, because relation names do not perfectly align, and objects in GPTKB are substantially more wordy. Third, we identify several properties not modelled at all in Wikidata, for instance, historical_significance (342K triples), art_style (84K triples), or hobbies (24K triples). This indicates that GPTKB potentially contains a significant amount of novel knowledge. A more comprehensive KB comparison in the style of Färber et al. (2018) is conceivable.